# Do Graph Neural Networks Dream of Landau Damping? Insights from Kinetic Simulations of a Plasma Sheet Model

### v2

@@ Abstract
Keywords: Plasma Physics, Machine Learning, Graph Neural Networks
@@ 1 Introduction
To obtain computational speed-ups, there have been recent attempts to combine existing PIC codes with machine learning surrogate models. These efforts include approaches to accelerate [5] or fully replace[6] the field solver block, or the integration of surrogate models into advanced physics extensions [7]. However, how to simultaneously obtain a significant computational gain, while enforcing known physics constraints is still an open research question.
@@
From the broad set of available surrogate models, one class of algorithms that can be of particular interest for kinetic plasma simulations are graph neural network-based approaches [19, 20], because of their capability to model both particle-particle [12] and particle-mesh interactions [13], as well as the possibility of enforcing known invariances or symmetries into the network architecture [20, 15, 16]. These approaches have been successfully applied to fluid [12, 13, 18], rigid body [13, 21], and charged particle dynamics [22, 15, 16]. However, to the best of our knowledge, their applicability to model kinetic plasma simulations is still to be demonstrated.
In this work we aim to model the predecessor of the PIC loop, the one-dimensional electrostatic sheet plasma model introduced by Dawson [23, 24]. This is an ideal initial testbench since it provides a simpler scenario, in terms of the problem structure and possible computational gains, while capturing a wide range of kinetic plasma physics phenomena that go beyond “collisionless” physics, including Coulomb collisions, and collisional thermalization [23, 25, 26, 27, 24]. Moreover, recent studies in the fundamental statistical physics processes in plasmas have been using the sheet model and/or direct extensions [28]. We show how to leverage previous work on graph neural network-based simulators by Sanchez-Gonzalez et al. [12] for kinetic plasma simulations and for the one-dimensional sheet model by introducing domain knowledge into the graph construction and simulator update mechanisms which enforce the desired symmetries and invariances. We discuss the advantages and disadvantages of using our surrogate model when compared to the standard physics simulator in terms of accuracy, run-time, energy conservation, and generalization capabilities. Finally, based on our findings, we comment on the expected impact of graph neural network-based simulators for the multi-dimensional PIC scenario.
@@ 3.1 Graph Representation
The plasma is represented as a graph with a set of nodes $\{\mathbf{n}_{i}\}$ representing the negatively charged sheets, and a set of directed edges $\{\mathbf{r}_{ij}\}$ which connect neighboring sheets. In our case, for simplicity and performance reasons we opted to connect only the first neighbors (additional comments on the impact of higher connectivity are provided throughout the text). Each node vector $\mathbf{n}_{i}$ contains the information relative to the corresponding negatively charged sheet, while each edge vector $\mathbf{r}_{ij}$ contains information about the relative displacement of node $i$ in relation to node $j$. They are defined as:
$\eqalign{\mathbf{n}^{t}_{i}&=\left[\xi^{t}_{i},v^{t}_{i}\right]\cr\mathbf{r}^{t}_{ij}&=\left[x^{t}_{j}-x^{t}_{i}\right]}$ (2)
@@
When considering reflecting boundary conditions, extra nodes representing mirrored versions of the boundary sheets are added to the graph (figure 4).
Figure 4: Graph representation of a four-sheet system for different boundary conditions. The number of mirror sheets used for reflecting boundaries is not necessarily one, and, if necessary, can be larger than the number of sheets inside the box.
Ideally, the number of mirror sheets should be large enough so that no boundary sheet crosses all mirror sheets in a single time step (in our case we set it to the number of message-passing steps, which we cover in Section 3.2). On the other hand, for periodic boundaries the graph becomes cyclic and we consider the distance between boundary sheets to be equal to the distance through the walls of the simulation box. In subsequent sections, we show that representing boundaries in this way allows the GNN model to be applied to different boundary conditions than the ones it was trained on, since it learns interactions between pairs of sheets (not interactions with the wall). However, this comes at the cost of an extra hard-coded step, which re-injects sheets that crossed the boundary (procedure to be explained in more detail in later sections).
We experimented with several other possible representations of the system, (e.g. different boundary representations, node and edge parameters, connect $n$-nearest neighbors) but these variants either produced worse results (in terms of accuracy or generalization capabilities) or introduced extra complexity and memory requirements that did not result in meaningful accuracy and/or speed-up improvements for the tested scenarios.
@@ 3.2 Graph Neural Network Architecture
The GNN module follows an encoder-processor-decoder architecture as introduced by Battaglia et al. [19]. A slight modification is used, which includes the sent edges in the node update mechanism of the processor block (equivalent to the Graph Network implementation available in Jraph [29]). The main building blocks are as follows:
Encoder: The encoder transforms each node $\mathbf{n}_{i}$ and edge $\mathbf{r}_{ij}$ vectors into a latent space node $\mathbf{v}_{i}$ and edge $\mathbf{e}_{ij}$ representation according to:
$\eqalign{\mathbf{v}_{i}=\varepsilon^{v}\left(\mathbf{n}_{i}\right)\cr\mathbf{e}_{ij}=\varepsilon^{e}\left(\mathbf{r}_{ij}\right)}$ (3)
where $\varepsilon^{v}$ and $\varepsilon^{e}$ are learnable functions.
Processor: The processor consists of a series of $M$ Graph Network (GN) blocks adapted from [19, 29], which sequentially update the node and edge values according to:
$\eqalign{\mathbf{e}_{ij}^{m+1}&=\phi^{e}\left(\mathbf{e}^{m}_{ij},\mathbf{v}^{m}_{i},\cdot\mathbf{v}^{m}_{j}\right)\cr\overline{\mathbf{e}}_{r_{i}}^{m+1}&=\sum_{j\in\mathcal{N}(i)}\mathbf{e}^{m+1}_{ij}\cr\overline{\mathbf{e}}_{s_{i}}^{m+1}&=\sum_{j\in\mathcal{N}(i)}\mathbf{e}^{m+1}_{ji}\cr\mathbf{v}_{i}^{m+1}&=\phi^{v}\left(\overline{\mathbf{e}}^{m+1}_{r_{i}},\overline{\mathbf{e}}^{m+1}_{s_{i}},\mathbf{v}^{m}_{i}\right)}$ (4)
@@
where the superscript denotes the block number, $\mathcal{N}(i)$ the set of nodes connected to $i$, and $\phi^{v}$, $\phi^{e}$ are learnable functions. The value of $M$ is set depending on the training time step and the maximum velocity of the sheets present in the training simulations. Ideally, $M$ should be larger than the maximum number of neighboring sheets that a given sheet crosses in any particular time step, since a graph node will at most receive/send information from/to the $M^{\text{th}}$ neighbor. This condition can be relaxed if the $N$-nearest neighboring nodes are directly connected (with the cost of additional memory requirements), and we have observed similar performance in our tested scenarios for models with an equivalent $M\!\times\!N$ factor. However, for higher crossing frequencies than the ones presented in this work, it might be preferable to increase $N$ to avoid possible information bottlenecks [30].
@@
Decoder: The decoder block transforms the latent node representations of the last layer into the output vector:
$\mathbf{y}_{i}=\delta^{v}\left(\mathbf{v}_{i}^{M}\right)$ (5)
@@
where $\delta^{v}$ is a learnable function. In our case, the output vector $\mathbf{y}_{i}$ is a single real value that corresponds to an estimate of the individual finite difference sheet acceleration $a^{t}_{i}=(v^{t+1}-v^{t})/\Delta t=(x^{t+1}-2x^{t}-x^{t-1})/\Delta t$ normalized to the intersheet spacing $\delta$.
@@
Although this GNN architecture, by design, does not enforce equivariance with respect to reflections (a symmetry present in the sheet model, i.e. if the simulation box is flipped the absolute value of the predicted accelerations should simply switch signs), we observed that the network was nonetheless capable of correctly approximating this symmetry within the training data range. In fact, we developed an alternative architecture that enforced this equivariance but did not observe relevant gains concerning the required number of training simulations nor improved rollout accuracy or energy conservation capabilities (in fact we observed a deterioration for considerable out-of-training data distribution values). Similarly, not using the sent messages for the node update mechanism led to poorer energy conservation for (considerably) out-of-training data distribution values. More details about these comparisons and the equivariant architecture are provided in D.
@@ 3.3 Position and Velocity Update
To update the sheet positions and velocities we use a semi-implicit first-order Euler integration scheme as follows:
$\eqalign{\tilde{v}^{t+1}&=v^{t}+a^{t}\Delta t\cr\tilde{x}^{t+1}&=x^{t}+\tilde{v}^{t+1}\Delta t\ .}$ (6)
@@ 5.4 Limitations
The first constraint arises because the network has to learn to predict sheet crossings, which implicitly forces it to know what is the simulation step. To train a single model for different simulation steps, it would be necessary to provide $\Delta t$ as an input to the network (in the GNS the time step only appears explicitly in the ODE integrator). Alternatively, as we have shown, different models can be trained, one per $\Delta t$. As long as enough training data is provided and the model architecture is scaled accordingly (e.g. by increasing the number of message passing steps for larger $\Delta t$ or connecting the $n^{\text{th}}$-closest neighbors) there is no limitation on the time step which can be used to train the model.
@@ 6 Recovering Known Kinetic Plasma Processes
In order to provide stronger evidence of the GNS generalization capabilities, we showcase a broad range of known kinetic plasma processes that the simulator is able to recover. These examples, present in both the original sheet model benchmarks [23, 25, 26, 27, 24] and other kinetic codes benchmarks [2, 36, 3, 28], aim to demonstrate the capability of the GNS to simulate collective behavior in accordance with known kinetic theory. An important point to stress is that the surrogate simulator was not explicitly trained to reproduce these effects. The GNN only learned to (correctly) model single-step updates over a reduced system size (10 sheets). However, when we apply it to larger system sizes and longer time durations, we observe the emergence of the expected kinetic collective plasma dynamics.

### v3

# Learning the dynamics of a one-dimensional plasma model with graph neural networks

@@ Abstract
Keywords: Plasma Physics, Kinetic Simulations, Machine Learning, Graph Neural Networks
@@ 1 Introduction
To obtain computational speed-ups, there have been recent attempts to combine existing PIC codes with machine learning surrogate models. These efforts include approaches to accelerate [5] or fully replace[6, 7] the field solver block, reduce the computational burden associated with the particle push and grid-particle/particle-grid interpolation [8, 9], and the integration of surrogate models into advanced physics extensions [10]. In parallel, PIC simulations and machine learning algorithms have also been used to train fast surrogate models for plasma accelerator setups [11, 12, 13, 14], to learn closures for fluid simulations [15], to model hybrid plasma representations [16], and to recover reduced plasma models [17]. However, obtaining a significant computational gain, while enforcing known physics constraints and reproducing the kinetic effects for a broad range of scenarios, is still an open research question.
@@
From the broad set of available surrogate models, one class of algorithms that can be of particular interest for kinetic plasma simulations are graph neural network-based approaches [29, 30, 31], because of their capability to model both particle-particle [22] and particle-mesh interactions [23], as well as the possibility of enforcing known invariances or symmetries into the network architecture [31, 25, 26]. These approaches have been successfully applied to fluid [22, 23, 28], rigid body [23, 32], and charged particle dynamics [33, 25, 26]. However, to the best of our knowledge, their applicability to model kinetic plasma simulations is still to be demonstrated. Graph-based representations and graph theory techniques have been explored in other branches of plasma physics, including low-temperature plasmas and plasma chemistry, but mostly for visualizing and reducing chemical reaction mechanisms [34, 35, 36].
In this work we aim to model the predecessor of the PIC loop, the one-dimensional electrostatic sheet plasma model introduced by Dawson [37, 38]. This is an ideal initial testbench since it provides a simpler scenario, in terms of the problem structure and possible computational gains, while capturing a wide range of kinetic plasma physics phenomena that go beyond “collisionless” physics, including Coulomb collisions, and collisional thermalization [37, 39, 40, 41, 38]. Moreover, recent studies in the fundamental statistical physics processes in plasmas have been using the sheet model and/or direct extensions [42]. We show how to leverage previous work on graph neural network-based simulators by Sanchez-Gonzalez et al. [22] for kinetic plasma simulations and for the one-dimensional sheet model by introducing domain knowledge into the graph construction and simulator update mechanisms which enforce the desired symmetries and invariances. We discuss the advantages and disadvantages of using our surrogate model when compared to the standard physics simulator in terms of accuracy, run-time, energy conservation, and generalization capabilities. In particular, we will show that when trained on data generated at high temporal resolution data (high-fidelity), the model is able to learn an improved algorithm to resolve sheet crossings (equivalent to resolving Coulomb collisions) for lower temporal resolutions when compared with the traditional physics solver. Futhermore, while the crossing correction algorithm implemented by the traditional solver is a serial routine, the model is also capable of reducing the overall run-time, since its operations are mostly parallelized. Finally, based on our findings, we comment on the expected impact of graph neural network-based simulators for the multi-dimensional PIC scenario.
@@ 3.1 Graph representation
The plasma is represented as a graph $\mathcal{G}$ with a set of nodes $\{\mathbf{n}_{i}\}$ representing the negatively charged sheets, and a set of directed edges $\{r_{ij}\}$ which connect neighboring sheets (figure 4). In our case, we opted to connect only the first neighbors (additional comments on the trade-off between higher connectivity and a deeper GNN are provided throughout the text). Each node $\mathbf{n}_{i}$ is represented by a vector containing the information relative to the corresponding negatively charged sheet, while each edge $r_{ij}$ contains the relative displacement of sheet $i$ in relation to sheet $j$. They are defined as:
$\eqalign{\mathbf{n}^{t}_{i}&=\left[\xi^{t}_{i},v^{t}_{i}\right]\cr r^{t}_{ij}&=x^{t}_{j}-x^{t}_{i}}$ (2)
@@
When considering reflecting boundary conditions, extra “guard” nodes representing mirrored versions of the sheets closer to the respective boundary are added to the graph (figure 4, mirrored versions follow $\xi^{m}_{i}=-\xi_{i}$, $v^{m}_{i}=-v_{i}$, for the left boundary $x^{m}_{eq_{i}}=x_{eq_{i}}-i\cdot\delta$, for the right boundary $x^{m}_{eq_{i}}=x_{eq_{i}}+(N+1-i)\cdot\delta$)).
Figure 4: Graph representation of a four-sheet system for different boundary conditions. Each graph node corresponds to a physical sheet. Directed edges connect neighboring nodes and represent the possible interaction between sheets. Computationally, nodes are represented by a vector and edges by a scalar according to equation (2). For the sheet model, boundaries are modeled using guard sheets [38]. For reflecting boundaries they represent mirrored versions of the sheet(s) closer to the boundary, for periodic boundaries they represent equal versions of the sheet(s) closer to the opposite wall. For the GNS, we adopt a similar representation for reflecting boundaries by adding guard nodes representing mirrored versions of the sheets closer to the boundary. The number of guard nodes is set to the number of message-passing steps of the GNN. For periodic boundaries, we instead add directed edges between the first and last node.
Ideally, the number of mirror sheets should be large enough so that no boundary sheet crosses all mirror sheets in a single time step (in our case we set it to the number of message-passing steps, which we cover in Section 3.2). On the other hand, for periodic boundaries, the graph becomes cyclic by introducing edges between the first and last nodes and considering the distance between the corresponding sheets to be equal to the distance through the walls of the simulation box ($r_{1N}=-r_{N1}=x_{N}-x_{1}-L$). In subsequent sections, we show that representing boundaries in this way allows the GNN model to be applied to different boundary conditions than the ones it was trained on, since it learns solely interactions between sheets (not interactions with the wall). However, this comes at the cost of an extra hard-coded step, which re-injects sheets that crossed the boundary (procedure to be explained in more detail in later sections).
We experimented with several other possible representations of the system, (e.g. different boundary representations, node and edge parameters) but these variants either produced worse results (in terms of accuracy or generalization capabilities) or introduced extra complexity and memory requirements that did not result in meaningful accuracy and/or speed-up improvements for the tested scenarios. Connecting $n$-nearest neighbors allowed us to achieve similar accuracies for shallower GNNs (since information propagates faster through the graph). However, it introduced extra memory requirements. It is therefore a trade-off that should be taken into account for future test scenarios.
@@ 3.2 Graph Neural Network architecture
The GNN module used (figure 5) follows an encoder-processor-decoder architecture as introduced by Battaglia et al. [30]. This architecture allows us to model the interactions between sheets by performing a series of message-passing steps between the connected nodes in the graph (representing the sheets) and therefore predict each individual sheet acceleration for the corresponding time step (which includes the contribution from the crossings and its oscillatory motion). The main building blocks are formalized as follows.
Figure 5: Schematic of the GNN encoder-processor-decoder architecture. The processor block is composed of $M$ independent Graph Network (GN) blocks similar to [30]. (a) The encoder block converts the input graph $\mathcal{G}^{t}$ node/edges ($\mathbf{n}_{i}^{t}$, $r_{ij}^{t}$) into a higher-dimension latent space graph representation $\mathbf{G}^{0}$ with node/edge vectors ($\mathbf{v}_{i}^{0}$, $\mathbf{e}_{ij}^{0}$). (b) Each $m$-th GN block updates the latent edge and node vectors of $\mathbf{G}^{m-1}\rightarrow\mathbf{G}^{m}$. The update functions $\phi^{m}_{e,v}$ are parameterized by a Multi-Layer Perceptron (MLP, in our case a 2-layer dense neural network). MLP weights are different for each GN block. (c) The decoder block computes the sheet acceleration $a_{i}^{t}$ from the corresponding final latent node representation $\mathbf{v}_{i}^{M}$. The linear transformations in the encoder/decoder block and the MLP weights from the GN blocks are learned during training.
Encoder: The encoder (figure 5a) transforms the graph $\mathcal{G}^{t}$ nodes $\mathbf{n}_{i}\in\mathbb{R}^{2}$ and edge $r_{ij}\in\mathbb{R}$ vectors into a (higher-dimensional) latent space graph representation $\mathbf{G}^{0}$, whose latent nodes $\mathbf{v}_{i}\in\mathbb{R}^{\mathrm{L}}$ and latent edges $\mathbf{e}_{ij}\in\mathbb{R}^{\mathrm{L}}$, are given by:
$\eqalign{\mathbf{v}_{i}=\varepsilon_{v}\left(\mathbf{n}_{i}\right)\cr\mathbf{e}_{ij}=\varepsilon_{e}\left(r_{ij}\right)}$ (3)
where $\varepsilon_{v}$ and $\varepsilon_{e}$ are learnable functions and L is the latent space size.
Processor: The processor consists of a series of $M$ Graph Network (GN) blocks (figure 5b) adapted from [30]. (These blocks were modified to include the sent edge information in the node update function, which is equivalent to the GN implementation available in Jraph [43].) Each $m$-th GN block updates the latent graph outputted by the previous layer $\mathrm{GN}^{m}(\mathbf{G}^{m-1})\rightarrow\mathbf{G}^{m}$ according to:
$\eqalign{\mathbf{e}_{ij}^{m}&=\phi^{m}_{e}\left(\mathbf{e}^{m-1}_{ij},\mathbf{v}^{m-1}_{i},\mathbf{v}^{m-1}_{j}\right)\cr\overline{\mathbf{e}}_{r_{i}}^{m}&=\sum_{j\in\mathcal{N}(i)}\mathbf{e}^{m}_{ij}\cr\overline{\mathbf{e}}_{s_{i}}^{m}&=\sum_{j\in\mathcal{N}(i)}\mathbf{e}^{m}_{ji}\cr\mathbf{v}_{i}^{m}&=\phi_{v}^{m}\left(\overline{\mathbf{e}}^{m}_{r_{i}},\overline{\mathbf{e}}^{m}_{s_{i}},\mathbf{v}^{m-1}_{i}\right)}$ (4)
@@
where the superscript denotes the block number, $\mathcal{N}(i)$ the set of nodes connected to $i$, and $\phi^{m}_{v}$, $\phi^{m}_{e}$ are learnable functions. The value of $M$ is set depending on the training time step and the maximum velocity of the sheets present in the training simulations. Ideally, $M$ should be larger than the maximum number of neighboring sheets that a given sheet crosses in any particular time step, since a graph node will at most receive/send information from/to the $M^{\text{th}}$ neighbor. This condition can be relaxed if the $N$-nearest neighboring nodes are directly connected (with the cost of additional memory requirements), and we have observed similar performance in our tested scenarios for models with an equivalent $M\!\times\!N$ factor. However, for higher crossing frequencies than the ones presented in this work, it might be preferable to increase $N$ to avoid possible information bottlenecks [44].
@@
Decoder: The decoder block (figure 5c) transforms the node vectors of the last latent graph $\mathbf{G}^{M}$ into a scalar:
$y_{i}=\delta_{v}\left(\mathbf{v}_{i}^{M}\right)\equiv a_{i}^{t}$ (5)
@@
where $\delta_{v}$ is a learnable function. In our case, the output $y_{i}$ is a single real value that corresponds to an estimate of the individual finite difference sheet acceleration $a^{t}_{i}=(v^{t+1}-v^{t})/\Delta t=(x^{t+1}-2x^{t}-x^{t-1})/\Delta t$ normalized to the intersheet spacing $\delta$.
@@
Although this GNN architecture, by design, does not enforce equivariance with respect to reflections over an equilibrium position (a symmetry present in the sheet model, i.e. if the simulation box is flipped the absolute value of the predicted accelerations should simply switch signs), we observed that the network was nonetheless capable of correctly approximating this symmetry within the training data range. In fact, we developed an alternative architecture that enforced this equivariance but did not observe relevant gains concerning the required number of training simulations nor improved rollout accuracy or energy conservation capabilities (in fact we observed a deterioration for considerable out-of-training data distribution values). Similarly, not using the sent messages for the node update mechanism led to poorer energy conservation for (considerably) out-of-training data distribution values. More details about these comparisons and the equivariant architecture are provided in D.
@@ 3.3 Position and velocity update
The predicted accelerations $\bi{a}^{t}$ provided by the GNN decoder are used to update the sheets dynamics. For this purpose, we use a semi-implicit first-order Euler integration scheme as follows:
$\eqalign{\tilde{v}_{i}^{t+1}&=v_{i}^{t}+a_{i}^{t}\Delta t\cr\tilde{x}_{i}^{t+1}&=x_{i}^{t}+\tilde{v}_{i}^{t+1}\Delta t}$ (6)
which corresponds to the “ODE Integrator” block in figure 3.
@@ 4.3 Training
Examples of the obtained training loss curves are depicted in figure 6. We observe a similar behavior across other models, with the validation loss stabilizing close to the last epoch. The learning rate scheduler is particularly helpful in the later stage of training to reduce the training loss oscillations across batches.
Figure 6: Examples of the training and validation loss evolution during training. For each epoch, $9900$ gradient updates are performed (1 per training simulation). The shaded area represents the minimum and maximum train loss across the batches for the corresponding epoch, mean value is presented in full line. Validation loss is computed at the end of the epoch over 100 held-out simulations. The results show that the training loss, on average, decreases monotonically and that the validation loss plateaus close to the defined epoch limit. No relevant gains were observed for longer training runs.
@@ 5.4 Limitations
The first constraint arises because the network has to learn to predict sheet crossings, which implicitly forces it to know what is the simulation step. To train a single model for different simulation steps, it would be necessary to provide $\Delta t$ as an input to the network (in the GNS the time step only appears explicitly in the ODE integrator). Alternatively, as we have shown, different models can be trained, one per $\Delta t$. As long as enough training data is provided and the model architecture is scaled accordingly (e.g. by increasing the number of message passing steps for larger $\Delta t$ or connecting the $n^{\text{th}}$-closest neighbors) there is no limitation on the time step which can be used to train the model. In contrast, the crossing correction algorithm in the SM requires that $\Delta t\leq\pi/2~{}\omega_{p}^{-1}$ (more details in A).
@@ 6 Recovering known kinetic plasma processes
In order to provide stronger evidence of the GNS generalization capabilities, we showcase a broad range of known kinetic plasma processes that the simulator is able to recover. These examples, present in both the original sheet model benchmarks [37, 39, 40, 41, 38] and other kinetic codes benchmarks [2, 50, 3, 9, 42], aim to demonstrate the capability of the GNS to simulate collective behavior in accordance with known kinetic theory. An important point to stress is that the surrogate simulator was not explicitly trained to reproduce these effects. The GNN only learned to (correctly) model single-step updates over a reduced system size (10 sheets). However, when we apply it to larger system sizes and longer time durations, we observe the emergence of the expected kinetic collective plasma dynamics.
@@ 6.3 Electrostatic fluctuations
6.4 Dispersion relation
The electrostatic waves propagating inside the plasma are expected to obey a particular dispersion relation, i.e. for a given angular wavenumber $k$ only a certain wave angular frequency $\omega$ is allowed. The ratio between the two quantities defines the wave phase velocity $v_{ph}=\omega/k$. For electrostatic waves, also known as Langmuir waves, the dispersion relation is given by [1]:
[…]
where $\hat{f}_{0}$ corresponds to the distribution function in velocity space. The solution will have both a real and imaginary part. The real part corresponds to the wave angular frequency, while the imaginary part corresponds to the inverse of the wave damping time (a phenomenon known as Landau damping which will be explored in Section 6.6).
To demonstrate that the GNS recovers the expected dispersion relation we perform a simulation of a thermal plasma with $\lambda_{D}=5~{}\delta$ and $N_{sheets}=10^{5}$ until $t_{max}=2\times 10^{3}~{}\omega_{p}^{-1}$. For each time step, we compute the electric field inside the box using a resolution of $\Delta x=1~{}\delta$ and then perform a 2D-FFT along the time and $x$-axis. The obtained power spectrum is presented in figure 15.
Figure 15: Dispersion relation obtained for the electrostatic waves propagating inside a thermal plasma. The GNS recovers the expected Langmuir waves dispersion relation and also reveals the presence of the non-collective ballistic (free-streaming) highly damped modes associated with single particle effects (phase velocity $v_{ph}$ corresponds to the sheet velocity). The ballistic modes can only exist for $v_{ph}\leq v_{max}$ (maximum sheet velocity). In this figure, they are mostly visible until $v_{ph}\approx 3.5v_{th}$ which encompasses $99.95\%$ of the sheet velocities across all time steps.
The signal corresponding to the Langmuir waves is clearly visible and in agreement with the numerical solution of the dispersion relation [53]. Additionally, it is also visible the presence of non-collective ballistic (free-streaming) modes [54, 55] associated with the individual particle motion ($\omega=kv_{sheet}$ where $v_{sheet}$ is the velocity of the sheet).